# Adaptive Control of Dubins Vehicle in the Presence of Loss of Effectiveness

### v2

@@ Abstract
The control of a Dubins Vehicle when subjected to a loss of control effectiveness in the turning rate is considered. A complex state-space representation is used to model the vehicle dynamics. An adaptive control design is proposed, with the underlying stability analysis guaranteeing closed-loop boundedness and tracking of a desired path. It is shown that a path constructed by waypoints and a minimum turn radius can be specified using a reference model which can be followed by the closed loop system. The control design utilizes the complex representation as well as a PID controller for the nominal closed-loop. How the design can be modified to ensure to ensure path following even in the presence input constraints is also discussed. Simulation studies are carried out to complement the theoretical derivations.
@@ I Introduction
This work leverages a complex representation of the DV proposed in [9] as opposed to the commonly used trigonometric representation. Such a representation is beneficial as the number of states are reduced by half, no approximation or linearization is required in the control design, speed and turning control loops can be decoupled, and rotations are handled by multiplication of complex exponentials. With this as a starting point, we propose and demonstrate an adaptive controller when parametric uncertainties are present. A nominal PID control design is chosen, which can guarantee closed-loop stability. This nominal design is replaced with adjustable parameters in the adaptive controller, whose parameters are updated by suitably leveraging the underlying complex state variables. Despite the state variables being complex, it is shown that a real positive definite Lyapunov function exists and that the norm of the tracking error goes to zero asymptotically. Furthermore, our results demonstrate that the proposed AC architecture outperforms a conventional PID structure in a LOE scenario.
@@
This paper is organized as follows. Background of the DV model, trigonometric and complex representations, and a compromised scenario involving a LOE are presented in Section II. In Section III, we introduce a reference model with outputs corresponding to a desired path. We then leverage the decoupling of speed and turning rate control loops to simplify the bilinear system into a linear system. In Section IV we derive a PID controller that leads to closed loop stability of the nominal DV. In the following Section V, an AC solution for the DV subject to LOE and input saturation is derived and analyzed. In Section VI we present numerical experiments that validate the complex domain AC solution and make a comparison the to PID controller. We conclude the paper in Section VII.
@@ V Adaptive Control Design
In this section, we show that an AC can be designed for the compromised DV vehicle (16) and that it leads to closed-loop stability and parameter learning. We first introduce an estimate $\hat{\lambda}$ in (28) and choose an adaptive control input as
$u_{2}=\frac{1}{\hat{\lambda}e^{i\psi}}\left[\frac{\mathbf{k}^{\top}\mathbf{e}}{iV_{a}}+u_{2ref}e^{i\psi_{ref}}\right].$ (33)
@@ V-B Adaptive Control Design with Input Constraints
where the second term $iV_{a}\hat{\lambda}\Delta_{sat}e^{i\psi}$ quantifies the effect due by control saturation. In order to minimize control clipping, we choose the magnitude of curvature $|\kappa_{ref}|$ (and correspondingly $\dot{\psi}_{ref}$) as
@@
$\dot{e}_{v}=\mathbf{k}^{\top}\mathbf{e}+\lambda\tilde{\theta}\Bigl{(}\mathbf{k}^{\top}\mathbf{e}+iu_{2\text{ref}}v_{\text{ref}}\Bigr{)}-\tilde{\lambda}\bigl{(}iV_{a}\Delta_{sat}e^{i\psi}\bigr{)},$ (52)
where $\tilde{\lambda}=\lambda-\hat{\lambda}$ denotes parameter of $\lambda$ errors. The additional term $-\tilde{\lambda}iV_{a}\Delta_{sat}e^{i\psi}$ reflects the effect of saturation on the velocity error. Following a similar procedure used to get (36), we obtain the following closed-loop error dynamics
@@
and ensure $\dot{V}\leq 0$. In effect, these update laws drive $\tilde{\theta}$ and $\tilde{\lambda}$ in a manner that compensates for both the parametric uncertainty and the extra “degradation” term caused by the saturated heading command. Similarly to (42), is easy to see that tracking error $\mathbf{e}\in L_{2}$. We then apply Barbalat’s lemma to conclude that $\mathbf{e}(t)\to 0$ as $t\rightarrow\infty$. It should be noted that we are able to determine the derivative of $V$ even though $\mathbf{e}\in\mathbb{C}$ due to the fact that $V$ can be expressed completely in terms of the real variables $\mathbf{e}_{\text{real}}$ and $\mathbf{e}_{\text{imag}}$.

### v3

@@ Abstract
The control of a Dubins Vehicle when subjected to a loss of control effectiveness in the turning rate is considered. A complex state-space representation is used to model the vehicle dynamics. An adaptive control design is proposed, with the underlying stability analysis guaranteeing closed-loop boundedness and tracking of a desired path. It is shown that a path constructed by waypoints and a minimum turn radius can be specified using a reference model which can be followed by the closed loop system. The control design utilizes the complex state-space representation as well as a PID controller for the nominal closed-loop. How the design can be modified to ensure path following even in the presence input constraints is also discussed. Simulation studies are carried out to complement the theoretical derivations.
@@ I Introduction
This work leverages a complex state-space representation of the DV proposed in [9] as opposed to the commonly used trigonometric representation. Such a representation is beneficial as the number of states are reduced by half, no approximation or linearization is required in the control design, speed and turning control loops can be decoupled, and rotations are handled by multiplication of complex exponentials. With this as a starting point, we propose and demonstrate an adaptive controller when parametric uncertainties are present. A nominal PID control design is chosen, which can guarantee closed-loop stability. This nominal design is replaced with adjustable parameters in the adaptive controller, whose parameters are updated by suitably leveraging the underlying complex state variables. Despite the state variables being complex, it is shown that a real positive definite Lyapunov function exists and that the norm of the tracking error goes to zero asymptotically. Furthermore, our results demonstrate that the proposed AC architecture outperforms a conventional PID structure in a LOE scenario.
@@
This paper is organized as follows. Background of the DV model, trigonometric and complex state-space representations, and a compromised scenario involving a LOE are presented in Section II. In Section III, we introduce a reference model with outputs corresponding to a desired path. We then leverage the decoupling of speed and turning rate control loops to simplify the bilinear system into a linear system. In Section IV we derive a PID controller that leads to closed loop stability of the nominal DV. In the following Section V, an AC solution for the DV subject to LOE and input saturation is derived and analyzed. In Section VI we present numerical experiments that validate the complex domain AC solution and make a comparison the to PID controller. We conclude the paper in Section VII.
@@ V Adaptive Control Design
In this section, we show that an AC can be designed for the compromised DV vehicle (16) and that it leads to closed-loop stability and parameter learning. We first introduce an estimate $\widehat{\theta}$ of $\frac{1}{\lambda}$ in (28) and choose an adaptive control input as
$u_{2}=\frac{{\widehat{\theta}}}{e^{i\psi}}\left[\frac{\mathbf{k}^{\top}\mathbf{e}}{iV_{a}}+u_{2ref}\,e^{i\psi_{ref}}\right].$ (33)
@@ V-B Adaptive Control Design with Input Constraints
where the second term $iV_{a}\widehat{\lambda}\Delta_{sat}e^{i\psi}$ quantifies the effect due to control saturation and $\widehat{\lambda}$ is an independent estimate of $\lambda$. In order to minimize control clipping, we choose the magnitude of curvature $|\kappa_{ref}|$ (and correspondingly $\dot{\psi}_{ref}$) as
@@
$\dot{e}_{v}=\mathbf{k}^{\top}\mathbf{e}+\lambda\tilde{\theta}\Bigl{(}\mathbf{k}^{\top}\mathbf{e}+iu_{2\text{ref}}v_{\text{ref}}\Bigr{)}-\tilde{\lambda}\bigl{(}iV_{a}\Delta_{sat}e^{i\psi}\bigr{)}$ (52)
$\text{where}\quad\tilde{\theta}=\widehat{\theta}-\theta,\quad\tilde{\lambda}=\lambda-\widehat{\lambda}.$
The additional term $-\tilde{\lambda}iV_{a}\Delta_{sat}e^{i\psi}$ reflects the effect of saturation on the velocity error. Following a similar procedure used to get (36), we obtain the following closed-loop error dynamics
@@
and ensure $\dot{V}\leq 0$. Equations (54) and (58) imply that all solutions of the adaptive system defined by equations (53) and (57) are globally bounded. It is also easy to see that $\mathbf{e}\in L_{2}$. By applying Barbalat’s lemma which states that if a uniformly continuous function has a finite $L_{2}$ norm, we can conclude that $\mathbf{e}(t)\to 0$ as $t\rightarrow\infty$. In summary, this establishes that the error $\mathbf{e}$ together with $r$, $v_{a}$, and $\psi$ remain bounded, and that $x_{d}$ tracks $x_{dref}$ asymptotically. It should be noted that (i) the adaptive laws in (57) adjust the control parameters $\widehat{\theta}$ and $\widehat{\lambda}$ in a way that compensates for both the parametric uncertainty and the extra “degradation” term caused by saturation; (ii) the derivative of $V$ is computable even though $\mathbf{e}\in\mathbb{C}$ due to the fact that $V$ can be expressed completely in terms of the real variables $\mathbf{e}_{\text{real}}$ and $\mathbf{e}_{\text{imag}}$.